OpenStudy (anonymous):

How do you use the distance formula to find lengths? I need an example please.

OpenStudy (sleepyjess):

lets say you want to find the distance between 2,3 and 7,10 distance=\(\sqrt{(x_2-x_1)+(y_2-y_1)}\) use 2,3 as \(x_1,~y_1\) and 7, 10 as \(x_2,~y_2\)

OpenStudy (sleepyjess):

whoops \(D=\sqrt{(x_2-x_1)^2+(y_1-y_2)^2}\) \(D=\sqrt{(7-2)^2+(10-3)^2}\) \(D=\sqrt{5^2+7^2}\) \(D=\sqrt{25+49}~~~~~~~~D=\sqrt{74}\)